OpenStudy (anonymous):

The volume of a cylinder increased by 8 times the original volume. How many times greater are each of the dimensions of that new cylinder? 2 3 4 8

OpenStudy (campbell_st):

they are only double... the reasoning... is that the linear measurements are cubed to get volume... test it with a cylinder radius 1 unit height 1 unit... find the volume then cylinder, radius 2 units and height 2 units then compare them

OpenStudy (anonymous):

@campbell_st , I got C)

OpenStudy (campbell_st):

well C is wrong... radius = 1 height = 1 \[V = \pi \times1^1 \times 1 = \pi\] radius = 2 and height = 2 \[V = \pi \times 2^2 \times 2 = 8\pi... \] if you triple the measurements the volume increased by 3^3 or 27 so r = 3 and h = 3 \[V = \pi \times 3^2 \times 3 = 27\pi\] quadruple the measurements the increase is 4^3 or 64 times hopefully it helps you to understand the concept

OpenStudy (campbell_st):

so is the volume has increase by 8 times take the cube root... \[\sqrt[3]{8} = 2\] so the linear measurements were doubled